# LongAudio-RAG: Event‑Grounded Question Answering over Multi‑Hour Long Audio

### v2

@@ Abstract
Long-duration audio is increasingly common in industrial and consumer settings, yet reviewing multi-hour recordings is impractical, motivating systems that answer natural‑language queries with precise temporal grounding and minimal hallucination. Existing audio–language models show promise, but long‑audio question answering remains difficult due to context‑length limits. We introduce LongAudio-RAG (LA-RAG), a hybrid framework that grounds Large Language Model (LLM) outputs in retrieved, timestamped acoustic event detections rather than raw audio. Multi‑hour streams are converted into structured event records stored in an SQL database, and at inference time the system resolves natural‑language time references, classifies intent, retrieves only the relevant events, and generates answers using this constrained evidence. To evaluate performance, we construct a synthetic long‑audio benchmark by concatenating recordings with preserved timestamps and generating template‑based question–answer pairs for detection, counting, and summarization tasks. Finally, we demonstrate the practicality of our approach by deploying it in a hybrid edge–cloud environment, where the audio grounding model runs on‑device on IoT‑class hardware while the LLM is hosted on a GPU‑backed server. This architecture enables low‑latency event extraction at the edge and high‑quality language reasoning in the cloud. Experiments show that structured, event‑level retrieval significantly improves accuracy compared to vanilla Retrieval-Augmented Generation (RAG) or text‑to‑SQL approaches.
@@
Vakada Naveen, Kartik Hegde, Arvind Krishna Sridhar, Yinyi Guo, Erik Visser Qualcomm Technologies Inc., nvakada@qti.qualcomm.com, karthegd@qti.qualcomm.com, arvisrid@qti.qualcomm.com, yinyig@qti.qualcomm.com, evisser@qti.qualcomm.com
@@ 1 Introduction
Retrieval-augmented generation (RAG) reduces hallucinations by grounding model outputs in retrieved evidence Lewis et al. (2020); Gupta et al. (2024). Although widely used for knowledge-intensive QA and summarization, RAG performance is tightly coupled to retrieval quality, especially for ambiguous or underspecified queries Karpukhin et al. (2020); Gupta et al. (2024). Recent work has extended these ideas beyond text: audio-integrated RAG retrieves directly from audio–text representations rather than relying solely on transcripts Chen et al. (2025), while long-context video understanding similarly benefits from retrieval-based selection of relevant segments to manage large input spaces and improve faithfulness Luo et al. (2024); Tan et al. (2025).
This work presents a hybrid framework for long audio question answering that grounds LLM outputs in timestamped acoustic detections rather than generating responses directly from multi-hour audio, for chatbot application as shown in Figure 1. The system first converts long audio into a sequence of detected events with event name, timestamps, confidence scores and related attributes, stored in a SQL database for efficient retrieval. At query time, temporal expressions are resolved into concrete intervals, intent is classified (e.g., detection, counting, summary), and only the relevant events are passed to the LLM for answer generation as shown in Figure 2.
To enable controlled evaluation of time-bounded queries missing in existing short-audio datasets Wijngaard et al. (2025); Gemmeke et al. (2017) we construct a synthetic long audio benchmark by concatenating industrial and external recordings while preserving event timestamps. Template-based question–answer pairs are generated across detection, counting, and summary tasks with diverse temporal expressions, allowing systematic testing of long audio reasoning.
Contributions.
We present a comprehensive solution to the open problem of performing reliable question answering over multi‑hour audio recordings that can span across days.
@@
    We introduce a hybrid grounding framework that anchors answers in timestamped acoustic events rather than raw audio, enabling precise temporal reasoning and reducing hallucination on hours-long content.
@@
    We proposed a complete, reproducible implementation stack, including event extraction, SQL-backed retrieval, temporal reference resolution, intent classification, and evidence-constrained generation.
@@
    We deploy the system in a hybrid edge-cloud configuration: audio grounding runs on IoT devices while the LLM is hosted on a GPU server, a split that reduces bandwidth usage, and scales language reasoning.
  • (iv)
    We design the system to be latency-aware by performing event detection and filtering at the edge, ensuring that only compact, relevant evidence is transmitted and enabling fast, stable responses even for multi-hour queries.
  • (v)
    We proposed a synthetic long audio benchmark and demonstrated improvements over text-to-SQL and RAG-based approaches.
To our knowledge, this is among the first event-level, edge–cloud systems for multi-hour audio question answering.
@@ 2 Related Work
Research on retrieval‑augmented generation (RAG) highlights the value of grounding LLM outputs in external information to reduce hallucinations and improve factuality. Canonical RAG pipelines pair dense retrievers with generators to condition responses on retrieved passages Lewis et al. (2020), with surveys documenting challenges such as retrieval ambiguity, domain shift, and bias in retrieval corpora Gupta et al. (2024). Dense retrievers like DPR enable semantic matching beyond token overlap, but often struggle when queries are underspecified Karpukhin et al. (2020), motivating hybrid retrieval strategies and query refinement. Extending RAG beyond text has gained traction in audio and video domains, where raw modalities exceed model context windows. Audio‑focused approaches range from ASR‑dependent pipelines to end‑to‑end systems such as WavRAG, which retrieves from audio–text hybrid knowledge without requiring transcription Chen et al. (2025). Strong pretrained encoders such as wav2vec 2.0 remain crucial for robust audio retrieval and reasoning Baevski et al. (2020). In videos, RAG-like methods retrieve auxiliary OCR/ASR/object metadata or select question‑relevant temporal segments, as in Video‑RAG Luo et al. (2024) and RAG‑Adapter Tan et al. (2025). Efficient video‑to‑text conversion frameworks like ViTA Arefeen et al. (2024) and spatiotemporal architectures including I3D Carreira and Zisserman (2017) and TimeSformer Bertasius et al. (2021) further support long‑video retrieval and understanding.
@@ 3.1 Automated QA generation
We generate evaluation QA pairs using a two-tiered approach that captures both controlled and realistic query variability for long‑audio understanding systems. Simple-QA pairs are created deterministically from ground‑truth event annotations produced by the Audio Grounding Model (AGM) module, covering detection, counting, and summary queries expressed in an unambiguous HH:MM:SS format. We created 3 types of queries - detection, counting, summary . To better reflect real‑world usage, we also produce complex-QA pairs that incorporate diverse temporal expression types including 24‑hour and 12‑hour formats,shift‑based references, before/after constraints, duration‑based phrasing, and other relative or segment-based intervals. More details about the QA pairs can be found in Appendix E.
@@ 3.2 Proposed approach for Long Audio Question Answering
First we process a long audio file by running our Audio Grounding Model (Section 3.2.1). using shifted windows to generate a text log of metadata ( Eg: audio event name , start time etc) as shown in Figure 2. This metadata is then inserted into a SQL database that can be queried. Then we retrieve the most relevant information from this database to answer the query. The detailed step by step process to generate the answering using the proposed method is given below:
@@ 3.2.1 Audio Grounding Model (AGM)
To preserve the flexibility required for on-device deployment across diverse acoustic environments and application scenarios, we develop an open-vocabulary Sound Event Detection (SED) model based on text to audio grounding. Instead of relying on a fixed label set, the model localizes sound events conditioned on free-form textual queries corresponding to the sounds of interest for a given use case. Our Audio Grounding Model (AGM) is trained following the phrase level WSTAG framework described in Xu et al. (2024), using the AudioCaps Kim et al. (2019) dataset from which we extract phrase-level supervision. AGM adopts the same architecture and hyperparameters as Xu et al. (2024): an audio encoder built from a CRNN with eight convolutional layers followed by a bidirectional GRU (BiGRU), and a text encoder composed of a single word embedding layer with mean pooling. Frame-level grounding scores are computed via the cosine similarity between audio frame embeddings and text tag embeddings, followed by a sigmoid activation. The output of this module is stored in JSON format which is then inserted into a SQLite database for question answering. The columns of the database are: audio event name, start time , end time , confidence of the generated class and other attributes like loudness.
@@ 3.2.2 Question Answering method
(4) SQL based retrieval and LLM based Response generation: After rephrasing, time resolution, and intent detection, the system retrieves relevant events from the database of audio events metadata. Broad queries use the full event set, whereas specific ones apply an embedding similarity based Tan et al. (2019) Top-k filter to use the most relevant audio events . An intent-specific prompt template (summary or detection/counting) then conditions the LLM on filtered, timestamped events to generate grounded and faithful responses. Anomaly queries follow a separate specialized workflow (Appendix I). More details on all these steps along with the prompts used can be found in Appendix F,G and M.
@@ 4 Results
We evaluate our system across multiple large language models (LLMs), data-processing pipelines, and query types to assess its effectiveness in real-world Industrial IoT audio analytics. The evaluation focuses on three downstream tasks derived from machine logs: event detection, event counting estimation, and natural language summarization.
@@ 4.1 Dataset for evaluation
We built a controlled synthetic benchmark using two 24-hour audio recordings, one for home IoT (HIoT) and one for industrial IoT (IIoT) by concatenating short, labeled events to preserve timestamps and vary density and noise (event classes listed in Appendix B). To match evaluation scope, we restrict acoustic classes to these predefined lists and pass the same closed set to downstream prompts, enabling grounded reasoning over multi-hour timelines without drift. Background audio uses industrial white-noise machine recordings for IIoT and low-hum ambience for HIoT, sourced from Freesound Fonseca et al. (2021). The final test vectors have a global SNR of 6 dB and are normalized to the Apple loudness target of –16 LUFS.
@@ 4.4 Audio Grounding Model evaluation
We evaluate the standalone AGM on long form audio using the Sound Event Detection (SED) task using DESED datasetSerizel et al. (2020) from the AudioMarathon benchmarkHe et al. (2025), a recent suite designed to assess long context audio understanding. For each multiple choice question (MCQ), we pass the four candidate sound event tags to AGM and compute a score for each tag by averaging the model’s frame level similarity scores across the entire audio clip. The predicted answer corresponds to the tag with the highest aggregated score. Under this simple inference setup, AGM achieves an F1 score of 74.8, ranking just below the top performing Qwen2.5 Omni 7B model, which attains 78.4 on the leaderboard ( Appendix D ). To obtain temporal event boundaries from AGM output, we apply a global decision threshold of 0.8 to the framewise scores and refine the resulting predictions using a 0.3 second median filter, which we hold constant across all experiments.
@@ 4.5 Comparison of Querying Approaches
To isolate the impact of high-quality event extraction, we compare against an AF3-based RAG pipeline. AF3 + RAG performs substantially worse, reinforcing that (i) well-structured AGM event logs are essential for downstream temporal reasoning, and (ii) retrieval grounded in event-level representations is more effective than operating on AF3 features or transcript-level content. We additionally conducted a human evaluation study using a curated set of 113 questions, divided into three categories: 67 detection, 35 counting, and 11 summarization questions. A human evaluator rated system responses on a 1–5 scale, on which our approach with Phi-3.5-MoE LLM, achieved an overall average score of 4.28. Further details on the human and GPT‑based evaluation setup and analysis are provided in Appendix H.
@@ 4.7 Time resolution module analysis
To evaluate our time‑resolution methods, we created 45 QA pairs spanning difficulty (easy, medium, hard) and semantic type (details in Appendix F). Easy questions use explicit, unambiguous expressions (e.g., “between 2pm and 4pm”), medium questions require interpreting relative phrases and simple arithmetic (e.g., “first 2 hours of day shift”), and hard questions involve ambiguous or colloquial references, boundary cases, or multi‑step constraints (e.g., “around lunchtime”). As shown in Table 3, the combined regex+LLM approach achieves the highest overall accuracy (77.78%), including perfect accuracy on easy cases and strong improvements on medium and hard queries.
@@ 4.8 Deployment Considerations
The system has been deployed in real-world environments with all components running in production. The Audio Grounding Model (AGM) operates on the Qualcomm IQ-9075 platform, which delivers up to 100 TOPS of on-device AI performance, enabling low-latency audio processing while keeping raw audio local for privacy. Its octa-core Kryo Gen6 CPU, Hexagon Tensor Processors, and industrial operating range support continuous edge inference. The LLM and embedding services run on NVIDIA A100 GPU servers, providing the throughput required for large-scale embedding generation and model inference. Other details of the software implementation can be found in Appendix K.
@@ 5 Conclusion
We presented an end-to-end framework for multi-hour audio understanding that integrates a lightweight AGM, structured event logs, and LLM-based reasoning. Our approach outperforms standard RAG and Text2SQL pipelines by leveraging temporally aligned AGM events for accurate detection, counting, and summarization. Experiments across diverse datasets show that mid-sized instruction-tuned and MoE models, when paired with high-quality event representations, deliver strong performance without relying on very large LLMs. Overall, the results demonstrate that structured event extraction combined with efficient retrieval forms a reliable foundation for machine-centric audio analytics. Future work includes expanding to additional sensing modalities and introducing agentic capabilities that leverage tools such as text-to-SQL and code generation. We also aim for fully end-to-end on-device deployment, running both the AGM and compact LLMs on edge hardware for low-latency inference.
Appendix A Ethical Considerations
This work involves processing long-duration audio, which may contain sensitive personal or environmental information. To mitigate privacy risks, all raw audio is processed locally on edge devices, and only structured event metadata is transmitted to the server, reducing the likelihood of exposing identifiable content. Our system does not attempt to infer personal attributes, emotions, or identities, and all event logs used for evaluation are synthetic or derived from publicly available datasets to avoid handling real user data. Finally, we ensure that model outputs are grounded in retrieved evidence to minimize hallucinations and reduce the risk of generating misleading or harmful responses.
Appendix B Audio classes for Home-IoT and Industrial IoT datasets
@@
Appendix C Custom sound enrollment and detection
@@ Appendix D AGM evaluation
Appendix E Automatic QA pair generation
E.1 Simple-QA pair generation
The development of robust evaluation frameworks for long audio based chatbot systems necessitates the creation of comprehensive question-answer (QA) datasets that can systematically assess model performance across diverse query types and temporal expressions. In this work, we present a simple yet effective QA generation methodology that employs a deterministic approach to create evaluation benchmarks for long-form audio understanding systems. Our framework generates three primary categories of queries: detection queries (binary presence/absence verification), counting queries (occurrence counting), and summary queries (temporal and statistical aggregation). To ensure temporal consistency and reproducibility, we adopt a standardized HH:MM:SS time format for all temporal expressions, eliminating ambiguities inherent in natural language time references such as "morning shift" or "12-hour format" variations. The generation process leverages ground-truth audio event annotations obtained using Audio Grounding Model (AGM) to construct queries with verifiable answers, incorporating both positive samples (events present in the audio) and negative samples (absent events and unrelated event categories) to evaluate false positive rates. Furthermore, we introduce synonym-based query variations to assess the semantic understanding capabilities of the target system, replacing canonical event labels with contextually equivalent terms. This simple generation approach produces balanced datasets with controlled distributions across query categories, enabling systematic evaluation of audio QA systems while maintaining interpretability and reproducibility, critical requirements for benchmarking in the audio understanding domain.
E.2 Complex-QA pair generation
While simple-QA generation provides a foundational evaluation framework, real-world audio understanding systems must handle the inherent complexity and variability of natural language temporal expressions. To address this challenge, we propose a complex-QA generation methodology that systematically incorporates diverse temporal reference formats to evaluate the robustness of audio event detection systems under realistic query conditions. Our approach extends beyond standardized HH:MM:SS formats to encompass seven distinct temporal expression types: (1) standard 24-hour format (e.g., "between 08:00 and 16:00"), (2) 12-hour format with meridiem indicators (e.g., "from 2:30 pm to 4:00 pm"), (3) shift-based references aligned with industrial or operational contexts (e.g., "during the morning shift"), (4) before/after temporal boundaries (e.g., "after 17:30"), (5) duration-based expressions relative to shifts or day boundaries (e.g., "first 2 hours of day shift"), (6) half-based temporal divisions (e.g., "second half of afternoon shift"), and (7) relative duration expressions (e.g., "between 1 hour and 2 hours"). By systematically varying temporal expression complexity, our QA generation facilitates the identification of model weaknesses in time parsing, semantic interpretation, and cross-format generalization, critical capabilities for deploying audio understanding systems in production environments where users employ diverse and unpredictable temporal references.
E.3 Implementation Details and Dataset Characteristics
Multi-Domain Support: The framework supports two configurable IoT domains: Industrial IoT (12 event classes: arc welder, stamping machine, diesel forklift, etc.) and Home IoT (13 event classes: baby cry, doorbell, glass breaking, etc.), each with domain-specific event taxonomies and synonym mappings to ensure contextual relevance.
Dataset Scale: The generation process produces 100 QA pairs per section across five generation phases: (1) detection & frequency with original event labels, (2) detection & frequency with synonym variations, (3) detection & frequency with unrelated events (negative samples), (4) specific event summaries, and (5) generic summaries. Since detection and frequency queries are generated in pairs for the first three sections, this yields approximately 800 total QA pairs (600 detection/frequency + 200 summary) with balanced positive/negative distributions and a minimum 10-second temporal window.
Temporal Expression Distribution: In complex-QA mode, temporal expressions are sampled using a probability-based strategy: 5% full-day queries, 15% shift-based (when enabled), 10% before/after boundaries, 10% duration-based, 5% half-based, 20% 12-hour format, and the remainder as standard 24-hour format, ensuring comprehensive coverage across temporal expression types.
Shift Definitions: For industrial contexts, standardized shifts are defined as: Morning/Day Shift (08:00–16:00), Afternoon/Evening Shift (16:00–24:00), and Night Shift (00:00–08:00). Shift-based queries can be toggled via a configuration flag, enabling evaluation with or without shift-specific temporal reasoning.
Ground Truth Structure: Each QA pair includes structured JSON metadata containing query category/subcategory, temporal bounds (HH:MM:SS), event tags, time expression type, and relevant statistics (detection status, occurrence counts, first/last event timestamps), enabling automated evaluation and error analysis.
Validation Mechanisms: The framework implements safeguards including configuration validation with automatic fallbacks, maximum attempt limits (10× target count) to prevent infinite loops, minimum duration enforcement (10 seconds), and automatic mode compatibility checks.
Appendix F Time resolution
The detailed category-level results in Appendix Table 5 highlight the specific categories that influence model performance. Categories with well-structured or explicitly defined temporal expressions such as explicit time ranges, shift-based cues, and full-day implicit references show consistently strong performance, particularly when using the combined approach. In contrast, categories involving greater variability or ambiguity, including relative durations, typos and variations, and edge cases, yield lower accuracy across all methods, reflecting their inherent complexity. This breakdown provides granular insight into error patterns and complements the difficulty-based analysis presented in the main paper. Detailed flowchart and prompt are shown in Figure 3 and 4.
Table 5: Category-wise accuracy results for the time resolution module.
Category Regex-only LLM-only Combined
explicit_time_ranges 100.00 100.00 100.00
shift_based 100.00 100.00 100.00
relative_durations 60.00 60.00 100.00
before_after 100.00 50.00 100.00
half_periods 100.00 50.00 100.00
full_day_implicit 100.00 100.00 100.00
typos_and_variations 27.27 72.73 72.73
edge_cases 20.00 20.00 30.00
Figure 3: Time resolution module
Figure 4: Prompt used in time resolution module
Appendix G Query rephrasing module
The detailed flowchart and the prompt used for query rephrasing module are shown in Figure 5 and 6.
Figure 5: Query Rephrasing Module
Figure 6: Prompt used in Query Rephrasing Module
Appendix H Dialogue evaluation using humans and GPT-4o as judge

### v3

@@ Abstract
Long-duration audio is increasingly common in industrial and consumer settings, yet reviewing multi-hour recordings is impractical, motivating systems that answer natural‑language queries with precise temporal grounding and minimal hallucination. Existing audio–language models show promise, but long‑audio question answering remains difficult due to context‑length limits. We introduce LongAudio-RAG (LA-RAG), a hybrid framework that grounds Large Language Model outputs in retrieved, timestamped acoustic event detections rather than raw audio. Multi‑hour streams are converted into structured event records stored in an SQL database, and at inference time the system resolves natural‑language time references, classifies intent, retrieves only the relevant events, and generates answers using this constrained evidence. To evaluate performance, we construct a synthetic long‑audio benchmark by concatenating recordings with preserved timestamps and generating template‑based question–answer pairs for detection, counting, and summarization tasks. Finally, we demonstrate the practicality of our approach by deploying it in a hybrid edge–cloud environment, where the audio grounding model runs on‑device on IoT‑class hardware while the LLM is hosted on a GPU‑backed server. This architecture enables low‑latency event extraction at the edge and high‑quality language reasoning in the cloud. Experiments show that structured, event‑level retrieval significantly improves accuracy compared to vanilla Retrieval-Augmented Generation (RAG) or text‑to‑SQL approaches.
@@
Naveen Vakada, Kartik Hegde, Arvind Krishna Sridhar, Yinyi Guo, Erik Visser Qualcomm Technologies Inc., {nvakada, karthegd, arvisrid, yinyig, evisser}@qti.qualcomm.com
@@ 1 Introduction
Retrieval-augmented generation reduces hallucinations by grounding model outputs in retrieved evidence Lewis et al. (2020); Gupta et al. (2024). Although widely used for knowledge-intensive QA and summarization, RAG performance is tightly coupled to retrieval quality, especially for ambiguous or underspecified queries Karpukhin et al. (2020); Gupta et al. (2024). Recent work has extended these ideas beyond text: audio-integrated RAG retrieves directly from audio–text representations rather than relying solely on transcripts Chen et al. (2025), while long-context video understanding similarly benefits from retrieval-based selection of relevant segments to manage large input spaces and improve faithfulness Luo et al. (2024); Tan et al. (2025).
Our contributions are:
@@
    We introduce an event-grounded framework for multi-hour audio question answering that anchors responses in timestamped acoustic events rather than raw audio, enabling precise temporal reasoning and reducing hallucinations on long-duration content.
@@
    We provide a complete and reproducible implementation stack—including event extraction, SQL-backed retrieval, temporal reference resolution, intent classification, and evidence-constrained generation—and deploy it in a hybrid edge-cloud configuration that reduces bandwidth usage while scaling language reasoning.
@@
    We design the system to be latency-aware through edge-side event detection and filtering, enabling fast and stable responses even for multi-hour recordings, and we propose a synthetic long-audio benchmark demonstrating improvements over text-to-SQL and RAG-based baselines.
To our knowledge, this is among the first event-level, edge-cloud systems for multi-hour audio question answering. We cannot provide a demo link because LongAudio‑RAG depends on a specialized IoT hardware setup that cannot be shared or hosted online. The system is not a standalone package—running it requires our team’s support, making a continuous online demo infeasible.
@@ 2 Related Work
Research on retrieval‑augmented generation highlights the value of grounding LLM outputs in external information to reduce hallucinations and improve factuality. Canonical RAG pipelines pair dense retrievers with generators to condition responses on retrieved passages Lewis et al. (2020), with surveys documenting challenges such as retrieval ambiguity, domain shift, and bias in retrieval corpora Gupta et al. (2024). Dense retrievers like DPR enable semantic matching beyond token overlap, but often struggle when queries are underspecified Karpukhin et al. (2020), motivating hybrid retrieval strategies and query refinement. Extending RAG beyond text has gained traction in audio and video domains, where raw modalities exceed model context windows. Audio‑focused approaches range from ASR‑dependent pipelines to end‑to‑end systems such as WavRAG, which retrieves from audio–text hybrid knowledge without requiring transcription Chen et al. (2025). Strong pretrained encoders such as wav2vec 2.0 remain crucial for robust audio retrieval and reasoning Baevski et al. (2020). In videos, RAG-like methods retrieve auxiliary OCR/ASR/object metadata or select question‑relevant temporal segments, as in Video‑RAG Luo et al. (2024) and RAG‑Adapter Tan et al. (2025). Efficient video‑to‑text conversion frameworks like ViTA Arefeen et al. (2024) and spatiotemporal architectures including I3D Carreira and Zisserman (2017) and TimeSformer Bertasius et al. (2021) further support long‑video retrieval and understanding.
@@ 3.1 Automated QA generation
We generate evaluation QA pairs using a two-tiered approach that captures both controlled and realistic query variability for long‑audio understanding systems. Simple-QA pairs are created deterministically from ground‑truth event annotations produced by the Audio Grounding Model (AGM) module, covering detection, counting, and summary queries expressed in an unambiguous HH:MM:SS format. We created three types of queries - detection, counting, summary . To better reflect real‑world usage, we also produce complex-QA pairs that incorporate diverse temporal expression types including 24‑hour and 12‑hour formats, shift‑based references, before/after constraints, duration‑based phrasing, and other relative or segment-based intervals.
@@ 3.2 Proposed approach for Long Audio Question Answering
First we process a long audio file by running our Audio Grounding Model (Section 3.2.1) using shifted windows to generate a text log of metadata (e.g., audio event name, start time etc) as shown in Figure 2. This metadata is then inserted into a SQL database that can be queried. Then we retrieve the most relevant information from this database to answer the query.
@@ 3.2.1 Audio Grounding Model (AGM)
To preserve the flexibility required for on-device deployment across diverse acoustic environments and application scenarios, we develop an open-vocabulary Sound Event Detection (SED) model based on text-to-audio grounding. Instead of relying on a fixed label set, the model localizes sound events conditioned on free-form textual queries corresponding to the sounds of interest for a given use case. Our Audio Grounding Model (AGM) is trained following the phrase level WSTAG framework described in Xu et al. (2024), using the AudioCaps Kim et al. (2019) dataset from which we extract phrase-level supervision. AGM adopts the same architecture and hyperparameters as Xu et al. (2024): an audio encoder built from a CRNN with eight convolutional layers followed by a bidirectional GRU (BiGRU), and a text encoder composed of a single word embedding layer with mean pooling. Frame-level grounding scores are computed via the cosine similarity between audio frame embeddings and text tag embeddings, followed by a sigmoid activation. The output of this module is stored in JSON format which is then inserted into a SQLite database for question answering. The columns of the database are: audio event name, start time, end time, confidence of the generated class and other attributes like loudness.
@@ 3.2.2 Question Answering method
(4) SQL based retrieval and LLM based Response generation: After rephrasing, time resolution, and intent detection, the system retrieves relevant events from the database of audio events metadata. Broad queries use the full event set, whereas specific ones apply an embedding similarity based Tan et al. (2019) Top-k filter to use the most relevant audio events. An intent-specific prompt template (summary or detection/counting) then conditions the LLM on filtered, timestamped events to generate grounded and faithful responses.
@@ 4 Results
We evaluate our system across multiple large language models (LLMs), data-processing pipelines, and query types to assess its effectiveness in real-world Industrial IoT audio analytics. The evaluation focuses on three downstream tasks derived from machine logs: event detection, event counting, and summarization.
@@ 4.1 Dataset for evaluation
We built a controlled synthetic benchmark using two 24-hour audio recordings, one for home IoT (HIoT) and one for industrial IoT (IIoT) by concatenating short, labeled events to preserve timestamps and vary density and noise (event classes in Appendix A). To match evaluation scope, we restrict acoustic classes to these predefined lists and pass the same closed set to downstream prompts, enabling grounded reasoning over multi-hour timelines without drift. Background audio uses industrial white-noise machine recordings for IIoT and low-hum ambience for HIoT, sourced from Freesound Fonseca et al. (2021). The final test vectors have a global SNR of 6 dB and are normalized to the Apple loudness target of -16 LUFS.
@@ 4.4 Audio Grounding Model evaluation
We evaluate the standalone AGM on long form audio using the Sound Event Detection (SED) task using DESED datasetSerizel et al. (2020) from the AudioMarathon benchmarkHe et al. (2025), a recent suite designed to assess long context audio understanding. For each multiple choice question (MCQ), we pass the four candidate sound event tags to AGM and compute a score for each tag by averaging the model’s frame level similarity scores across the entire audio clip. The predicted answer corresponds to the tag with the highest aggregated score. Under this simple inference setup, AGM achieves an F1 score of 74.8, ranking just below the top performing Qwen2.5 Omni 7B model, which attains 78.4 on the leaderboard ( Appendix C ). To obtain temporal event boundaries from AGM output, we apply a global decision threshold of 0.8 to the framewise scores and refine the resulting predictions using a 0.3 second median filter, which we hold constant across all experiments.
@@ 4.5 Comparison of Querying Approaches
To isolate the impact of high-quality event extraction, we compare against an AF3-based RAG pipeline. AF3 + RAG performs substantially worse, reinforcing that (i) well-structured AGM event logs are essential for downstream temporal reasoning, and (ii) retrieval grounded in event-level representations is more effective than operating on AF3 features or transcript-level content. We additionally conducted a human evaluation study using randomly collected set of 113 question-response pairs, divided into three categories: 67 detection, 35 counting, and 11 summarization questions. An expert evaluator with experience in audio and industrial settings rated the system outputs on a 1–5 scale, on which our approach with Phi-3.5-MoE LLM, achieved an overall average score of 4.28. (Further details in Appendix D).
@@ 4.7 Time resolution module analysis
To evaluate our time‑resolution methods, we created 45 QA pairs spanning difficulty (easy, medium, hard) and semantic type. Easy questions use explicit, unambiguous expressions (e.g., “between 2pm and 4pm”), medium questions require interpreting relative phrases and simple arithmetic (e.g., “first 2 hours of day shift”), and hard questions involve ambiguous or colloquial references, boundary cases, or multi‑step constraints (e.g., “around lunchtime”). As shown in Table 3, the combined regex+LLM approach achieves the highest overall accuracy (77.78%), including perfect accuracy on easy cases and strong improvements on medium and hard queries.
@@ 4.8 Deployment Considerations
The system has been deployed in real-world environments with all components running in production. The Audio Grounding Model (AGM) operates on the Qualcomm IQ-9075 platform, which delivers up to 100 TOPS of on-device AI performance, enabling low-latency audio processing while keeping raw audio local for privacy. Its octa-core Kryo Gen6 CPU, Hexagon Tensor Processors, and industrial operating range support continuous edge inference. The LLM and embedding services run on NVIDIA A100 GPU servers, providing the throughput required for large-scale embedding generation and model inference. This is currently a restricted, proprietary system; access beyond the demo is available under a commercial license. More system details can be found in Appendix F and G.
@@ 5 Conclusion
We presented an end-to-end system for multi-hour audio understanding that integrates a lightweight AGM, structured event logs, and LLM-based reasoning. Our approach outperforms standard RAG and Text2SQL pipelines by leveraging temporally aligned AGM events for accurate detection, counting, and summarization. Experiments across diverse datasets show that mid-sized instruction-tuned and MoE models, when paired with high-quality event representations, deliver strong performance without relying on very large LLMs. Overall, the results demonstrate that structured event extraction combined with efficient retrieval forms a reliable foundation for machine-centric audio analytics. Future work includes adding more sensing modalities and agentic capabilities such as text-to-SQL and code generation, as well as achieving fully end-to-end on-device deployment with both the AGM and compact LLMs running on edge hardware for low-latency inference.
Appendix A Audio classes for Home-IoT and Industrial IoT datasets
@@
Appendix B Custom sound enrollment and detection
@@ Appendix C AGM evaluation
Figure 3: Software Implementation Architecture
Figure 4: LongAudio-RAG user interface
Appendix D Dialogue evaluation using humans and GPT-4o as judge